OpenStudy (anonymous):

When you solve an equation, how do you know when to add and when to subtract? You may include an example if you need to.

OpenStudy (anonymous):

Depends which one you need to express the unknown (x mostly) Example x-23=1 You need to add 23 so that x-23+23=1+23 x=24

OpenStudy (anonymous):

\[\Huge \frak Andras The Wizard\]

OpenStudy (anonymous):

:D Thanks